News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 10 September 1999 10.00am EST

During the last 24 hours the on-board load continued to operate as expected with HRC-S/LETG focus plate measurements on Capella and HRC-I observations of the science target CEN A. Preparations continue for the next load scheduled to begin late on 9/11.

Analysis of ACIS calibration source data from the last 5 days has shown an unexplained degredation in the energy resolution of the front-side illuminated CCD chips of ACIS. The degredation is evident in data starting from 5 days ago and shows a change in the FWHM from approx 130 eV to 500 eV. A series of diagnostic measurements were scheduled during the real-time passes yesterday and further diagnostic work will continue over the next 24 hours in order to understand the behavior. The ACIS engineering team is working to develop a fault tree to ensure all possible causes are covered.





Return to Status Reports